OpenStudy (anonymous):

Evaluate the expression below at x = 9. A. 6,606 B. 8 C. 126 D. 86

OpenStudy (anonymous):

https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{45\%20+\%20x^{3}}{x}

OpenStudy (greatlife44):

cant see your equation

OpenStudy (anonymous):



OpenStudy (anonymous):

@563blackghost

OpenStudy (ace-n-it):

k

OpenStudy (anonymous):

HELP ANYONE

OpenStudy (anonymous):

HELP ANYONE

OpenStudy (ace-n-it):

one sec

563blackghost (563blackghost):

We know that x is 9 so we input.... \(\Huge{\frac{45+9(3)}{9}}\) Simplify... \(\Huge\color{purple}{\frac{45+27}{9}}\) Add.... \(\Huge\color{red}{\frac{72}{9}}\) So what would the answer be?

OpenStudy (mathmath333):

which is ur question ? \(\large \color{black}{\begin{align} & \dfrac{45+x^{3}}{x}\hspace{.33em}\\~\\ \end{align}}\) or \(\large \color{black}{\begin{align} & \dfrac{45+x3}{x}\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

the second one

OpenStudy (ace-n-it):

ok

OpenStudy (ace-n-it):

i know it

OpenStudy (anonymous):

its 8 right

563blackghost (563blackghost):

We know that x equals 9 in which we can input into the equation as x....

OpenStudy (ace-n-it):

yup

563blackghost (563blackghost):

Yup ^^

OpenStudy (anonymous):

thxs bye

OpenStudy (ace-n-it):

np

563blackghost (563blackghost):

np ^^ byeeeee